OpenStudy (anonymous):

evaluate cos(arcsin(3/5) - arccos(3/5))

OpenStudy (anonymous):

Use the double angle formula for cosine and definition of inverses. So,

OpenStudy (anonymous):

\[\cos (\sin^{-1}\frac{3}{5}-\cos^{-1}\frac{3}{5})\]\[=\cos (\sin^{-1}\frac{3}{5}) \cos(\cos^{-1}\frac{3}{5})+\sin(\sin^{-1}\frac{3}{5})\sin(\cos^{-1}\frac{3}{5})\]

OpenStudy (anonymous):

\[=\sqrt{1-\sin^2(\sin^{-1}\frac{3}{5}})\frac{3}{5}+\frac{3}{5}\sqrt{1-\sin^2(\sin^{-1}\frac{3}{5})}\]

OpenStudy (anonymous):

\[=\sqrt{1-\left( \frac{3}{5} \right)^2}\frac{3}{5}+\frac{3}{5}\sqrt{1-\left( \frac{3}{5} \right)^2}\]\[=2.\frac{3}{5}\sqrt{\frac{16}{25}}=\frac{6}{5}\frac{4}{5}=\frac{24}{25}\]

OpenStudy (anonymous):

Fin.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

No probs. Become a fan ;)

OpenStudy (anonymous):

ok i will

OpenStudy (anonymous):

Thanks